OpenStudy (anonymous):

What value is added to both sides of the equation x2 – 4x = –2 in order to solve by completing the square? (1 point) 4 2 –2 –4

OpenStudy (mathstudent55):

Take half of the coefficient of the x term, and square it.

OpenStudy (campbell_st):

in a quadratic \[ax^2 + bx = c\] you need to add \[(\frac{b}{2})^2\] to both sides

OpenStudy (anonymous):

so i add 2 to both sides?

OpenStudy (mathstudent55):

Half of -4 is -2. Square of -2 is 4. Add 4 to both sides.

OpenStudy (campbell_st):

nope.... read the information again

OpenStudy (anonymous):

x2 – 4x = –2 add two in both sides so you get the quadratic equation and simplify that

OpenStudy (amoodarya):